OpenStudy (anonymous):

Write the standard form of the equation of the circle with the given characteristics center (-7,-4) radius 7

OpenStudy (tkhunny):

This is an eyeball problem. You should write it down about as quickly as you can read the problem statement. Standard form of a circle is (x-h)^2 + (y-k)^2 = r^2, where the center is (h,k) and the radius is r. Since this is EXACTLY the information we have been given, we are done! (x-(-7))^2 + (y-(-4))^2 = 7^2 This can be simplified a little: (x+7)^2 + (y+4)^2 = 49 Seriously, get good at it and get fast!

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."